Calculate the moles of C2H6O in 5.66x1024 molecules of C2H6O. Calculate each of the following quantities in 1.50 moles H3PO4: moles of H, moles of O, atoms of P, atoms of O
Answer #1

The solution is given below.

No. of particles Number of moles = 6:02 x 1023 No. of particles = 5.66 * 184 No. of males of 2400 = 5.66 *1024 6.02 X1023 = 0